OpenStudy (partypoison):

find the coordinates of the midpoints of a segment having endpoints Q(-0.2, 9.4) and R(2.5, -0.8)

OpenStudy (kenrick143):

Mid point equation x=1/2(xsub2-xsub1) y=1/2(ysub2-ysub1)

OpenStudy (wckeller):

Midpoint formula: \[x=(x_2-x_1)/2\]\[y=(y_2-y_1)/2\] x=(2.5-(-0.2))/2 y=((-0.8)-9.4)/2 Midpoint of the segment = (1.35,-5.1)